OpenStudy (anonymous):

For the function y=3x^2-5x+7x^-1, find a simple basic function right end behavior model?

OpenStudy (anonymous):

y = 3x^2 - 5x

OpenStudy (anonymous):

how did you get that answer? thanks!

OpenStudy (anonymous):

as x tends to infinity, 7x^-1 tends to 0. So the function tends to 3x^2 - 5x

OpenStudy (anonymous):

Okay what if the question asks: a simple basic function left end behavior model?

OpenStudy (anonymous):

same function as right end behavior

OpenStudy (anonymous):

so are the functions you wrote the left and right end behavior functions?

OpenStudy (amoodarya):

for large amount of x , they have almost same behavior

OpenStudy (amoodarya):

now i give you an example it clearly shows the same behavior of y=x^2 and y=x^2+1/x for large x (end behavior)

OpenStudy (anonymous):

oohh okay thank you soo much!